# Improved Approximation Algorithms for Chromatic and Pseudometric-Weighted Correlation Clustering

### v1

@@ Abstract
Correlation Clustering (CC) is a foundational problem in unsupervised learning that models binary similarity relations using labeled graphs. While classical CC has been widely studied, many real-world applications involve more nuanced relationships, either multi-class categorical interactions or varying confidence levels in edge labels. To address these, two natural generalizations have been proposed: Chromatic Correlation Clustering (CCC), which assigns semantic colors to edge labels, and pseudometric-weighted CC, which allows edge weights satisfying the triangle inequality. In this paper, we develop improved approximation algorithms for both settings. Our approach leverages LP-based pivoting techniques combined with problem-specific rounding functions. For the pseudometric-weighted correlation clustering problem, we present a tight $\frac{10}{3}$-approximation algorithm, matching the best possible bound achievable within the framework of standard LP relaxation combined with specialized rounding. For the Chromatic Correlation Clustering (CCC) problem, we improve the approximation ratio from the previous best of $2.5$ to $2.15$, and we establish a lower bound of $2.11$ within the same analytical framework, highlighting the near-optimality of our result.
@@ 1 Introduction
Clustering is a fundamental task in unsupervised learning, where the goal is to partition a set of objects into groups based on their pairwise relationships. One prominent problem in this domain is Correlation Clustering (CC) bansal2004correlation , which models binary similarity/dissimilarity between items using an edge-labeled graph: similar pairs are marked with a ‘$+$’ label and dissimilar pairs with a ‘--’. The objective is to partition the nodes to minimize disagreements—i.e., cases where the partitioning contradicts the edge labels. Due to its flexibility in not requiring a predefined number of clusters, CC has been widely utilized in various areas such as detecting communities in networks Chen2012 , inferring labels from user interactions Agrawal2009 ; Chakrabarti2008 and resolving ambiguous entities Kalashnikov2008 .
@@ 1.1 Related Works
The Correlation Clustering (CC) problem has been widely studied since its introduction ben-dor1999clustering , and it is known to be APX-hard, leading to efforts to develop approximation algorithms. Early work by Bansal, Blum, and Chawla introduced a constant-factor approximation algorithm bansal2004correlation . Charikar et al. CHARIKAR2005360 improved this to a 4-approximation using linear programming. Ailon, Charikar, and Newman then introduced the Pivot algorithm ailon2008aggregating , which achieved a 3-approximation in linear time. Chawla et al. chawla2015near further improved this to 2.06 using more refined LP-rounding techniques. More recently, researchers have surpassed the 2-approximation barrier. Cohen-Addad, Lee, and Newman cohenaddad2022sherali used the Sherali-Adams hierarchy to develop a $(1.994+\varepsilon)$-approximation, while Cohen-Addad et al. cohenaddad2023preclustering proposed preclustering, which improved the approximation to $(1.73+\varepsilon)$. The most recent breakthrough by Cao et al. cao2024clusterlp introduced the cluster LP, which unifies all known LP relaxations for CC. They show that this can be approximated efficiently using preclustering, achieving a $(1.437+\varepsilon)$-approximation, the best known guarantee for CC so far. In a more recent work cao2024clusterlp , they introduced a new approach to find a feasible solution for the cluster LP in sublinear time.
Chromatic Correlation Clustering (CCC) is an extension of the classical Correlation Clustering (CC) problem, where edge colors represent different types of relationships. Bonchi et al. bonchi2012chromatic introduced CCC and proposed a heuristic algorithm without approximation guarantees. Anava et al. anava2015chromatic presented a 4-approximation algorithm based on LP rounding and also proposed two practical methods: Reduce and Cluster (RC), with an approximation ratio of 11, and a heuristic named Deep Cluster (DC). Klodt et al. klodt2021color showed that the Pivot algorithm by Ailon et al. ailon2008aggregating , originally designed for standard CC, provides a 3-approximation guarantee for CCC. They also established that the RC algorithm achieves a 5-approximation for CCC. However, Pivot’s color-agnostic nature limits its performance in practice. To address this, Klodt et al. proposed a heuristic called Greedy Expansion (GE), which performs well empirically despite lacking theoretical guarantees. More recently, Xiu et al. XiuHTCH22 developed a 2.5-approximation algorithm for CCC based on a linear programming approach, improving upon the previous best-known ratio. They also introduced a greedy heuristic that achieves strong empirical results.
@@ 1.2 Our Results
  • For the Chromatic Correlation Clustering (CCC) problem, we enhance the LP-based method through a new analysis that yields a $2.15$-approximation, improving upon the previous best bound of $2.5$ by Xiu et al. XiuHTCH22 . We also establish a lower bound of $2.11$ within the same analytical framework, underscoring the near-optimality of our approach.
@@ 2 Preliminaries
The correlation clustering (CC) problem takes as input a signed undirected graph $G=(V,E=E^{+}\uplus E^{-})$, where each edge $e=uv\in E$ is assigned a sign ‘$+$’ or ‘--’, described by $e\in E^{+}$ or $e\in E^{-}$. The objective is to find a partition of the nodes such that the number of disagreements—i.e., negative edges within the same cluster and positive edges between different clusters—is minimized. In other words, the cost of the clustering $\mathcal{C}$ is as follows:
@@ 2.1 Pseudometric-weighted Correlation Clustering
The weighted Correlation Clustering (CC) problem is a generalization of the classical CC problem in which each edge is associated with a nonnegative violation cost. Specifically, for each edge $uv$, a weight $w_{uv}\geq 0$ is provided, and violating the edge’s label (either ‘$+$’ or ‘--’) incurs a penalty of $w_{uv}$. This differs from the standard setting, where all violations incur a uniform cost of $1$. The weighted variant allows us to encode edge-wise reliability: when $w_{uv}$ is large, it is more reasonable to follow the given label between $u$ and $v$.
However, assuming the Unique Games Conjecture, no $O(1)$-approximation algorithm exists for the general weighted case DEMAINE2006172 . An exception occurs when the weight function satisfies the triangle inequality, i.e., the weights form a pseudometric. In this pseudometric-weighted setting, a constant-factor approximation is known DBLP:conf/soda/CharikarG24 . Following the analysis of Charikar and Gao with $L=2$ yields an approximation factor of $\overline{B}_{\text{HR}}+\frac{1}{\frac{1}{3}}\leq\frac{4}{\frac{1}{3}}+2(L-1)+\frac{1}{\frac{1}{3}}=17,$ since the second type of charge occurs at most $L-1=1$ time in the charging scheme. The following is a natural LP relaxation of the weighted CC problem, extending (CC-LP):
@@ 2.2 Chromatic Correlation Clustering Problem
The Chromatic Correlation Clustering (CCC) problem is a variant of the classical CC problem in which each cluster is additionally assigned a color bonchi2012chromatic . The input includes a set of $L$ possible colors, as well as a special color $\gamma$ that denotes that two vertices should not be placed in the same cluster—analogous to a negative (‘--’) edge in the classical CC setting. When $L=1$ (i.e., a single cluster color), CCC reduces to the standard CC problem.
@@
Since CCC generalizes the standard CC problem, the integrality gap of (CCC-LP) is at least as large as that of (CC-LP), which is 2.
@@ 3 Approximation Algorithm
Input: Graph $G=(V,\,E)$, color function $\phi:E\rightarrow L\cup\{\gamma\}$, LP solution $\{x_{u}^{c}\}_{u\in V,\,c\in L}$ and $\{x_{uv}^{c}\}_{uv\in E,\,c\in L}$.
@@
     if $\exists c\in L$ s.t. $x_{u}^{c}<\frac{1}{2}$then
@@
     and $E^{+}\uplus E^{-}\uplus E^{\circ}$ is defined as a partition by color $c,\,\gamma,L\backslash\{c\}$ respectively.
@@ 4 Rounding Functions
The effectiveness of the LP-Pivot and LP-CCC algorithms critically depends on the choice of rounding functions used in the clustering process. Rounding functions $f^{+},\,f^{-},\,f^{\circ}:[0,1]\rightarrow[0,1]$ convert the LP value $x_{uv}$ to the non-selection probability $p_{uv}$ chawla2015near . The sign of the edge $uv$—either ‘$+$’, ‘--’, or $\circ$’—determines which rounding function is applied. The sign $\circ$’ indicates that the edge does not belong to $E$. The following natural conditions are imposed on any rounding function $f$:
@@ 4.3 Comparison with Prior Work
Pseudometric-weighted CC: The LP-UMVD-Pivot algorithm, recently proposed by Charikar and Gao DBLP:conf/soda/CharikarG24 for the Ultrametric Violation Distance (UMVD) problem, follows a pivoting-based rounding strategy applied to an LP relaxation. When the number of distinct pairwise distances between elements, denoted by $L$, is equal to 2, their algorithm can be viewed as a special case of our LP-Pivot framework. In this setting, the two distances $d_{1}$ and $d_{2}$ correspond to the ‘--’ and ‘$+$’ labels, respectively, used in our rounding procedure. The rounding functions used are:
[…]
@@
With triple-based analysis, this choice yields an approximation factor of $6$, as shown in the Appendix.

### v2

@@ Abstract
Correlation Clustering (CC) is a foundational problem in unsupervised learning that models binary similarity relations using labeled graphs. While classical CC has been widely studied, many real-world applications involve more nuanced relationships, either multi-class categorical interactions or varying confidence levels in edge labels. To address these, two natural generalizations have been proposed: Chromatic Correlation Clustering, which assigns semantic colors to edge labels, and pseudometric-weighted Correlation Clustering, which allows edge weights satisfying the triangle inequality. In this paper, we develop improved approximation algorithms for both settings. Our approach leverages LP-based pivoting techniques combined with problem-specific rounding functions. For the pseudometric-weighted correlation clustering problem, we present a tight $\frac{10}{3}$-approximation algorithm, matching the best possible bound achievable within the framework of standard LP relaxation combined with specialized rounding. For the Chromatic Correlation Clustering (CCC) problem, we improve the approximation ratio from the previous best of $2.5$ to $2.15$, and we establish a lower bound of $2.11$ within the same analytical framework, highlighting the near-optimality of our result.
@@ 1 Introduction
Clustering is a fundamental task in unsupervised learning, where the goal is to partition a set of objects into groups based on their pairwise relationships. One prominent problem in this domain is Correlation Clustering (CC) bansal2004correlation , which models binary similarity/dissimilarity between items using an edge-labeled graph: similar pairs are marked with a ‘++’ label and dissimilar pairs with a ‘-’. The objective is to partition the nodes to minimize disagreements—i.e., cases where the partitioning contradicts the edge labels. Due to its flexibility in not requiring a predefined number of clusters, CC has been widely utilized in various areas such as detecting communities in networks Chen2012 , inferring labels from user interactions Agrawal2009 ; Chakrabarti2008 and resolving ambiguous entities Kalashnikov2008 .
@@ 1.1 Related Works
The Correlation Clustering problem has been widely studied since its introduction ben-dor1999clustering , and it is known to be APX-hard, leading to efforts to develop approximation algorithms. Early work by Bansal, Blum, and Chawla introduced a constant-factor approximation algorithm bansal2004correlation . Charikar et al. CHARIKAR2005360 improved this to a 4-approximation using linear programming. Ailon, Charikar, and Newman then introduced the Pivot algorithm ailon2008aggregating , which achieved a 3-approximation in linear time. Chawla et al. chawla2015near further improved this to 2.06 using more refined LP-rounding techniques. More recently, researchers have surpassed the 2-approximation barrier. Cohen-Addad, Lee, and Newman cohenaddad2022sherali used the Sherali-Adams hierarchy to develop a $(1.994+\varepsilon)$-approximation, while Cohen-Addad et al. cohenaddad2023preclustering proposed preclustering, which improved the approximation to $(1.73+\varepsilon)$. The most recent breakthrough by Cao et al. cao2024clusterlp introduced the cluster LP, which unifies all known LP relaxations for CC. They show that this can be approximated efficiently using preclustering, achieving a $(1.437+\varepsilon)$-approximation, the best known guarantee for CC so far. In a more recent work cao2024clusterlp , they introduced a new approach to find a feasible solution for the cluster LP in sublinear time.
Chromatic Correlation Clustering is an extension of the classical Correlation Clustering problem, where edge colors represent different types of relationships. Bonchi et al. bonchi2012chromatic introduced CCC with a heuristic lacking guarantees. Anava et al. anava2015chromatic gave a 4-approximation via LP rounding, plus two practical methods: Reduce and Cluster (RC, ratio 11) and Deep Cluster (DC). Klodt et al. klodt2021color showed that Pivot ailon2008aggregating yields a 3-approximation and that RC achieves 5, but Pivot’s color-agnostic design limits performance; they instead proposed Greedy Expansion (GE), a heuristic effective in practice without guarantees. More recently, Xiu et al. XiuHTCH22 developed a 2.5-approximation algorithm for CCC based on a linear programming approach, improving upon the previous best-known ratio. They also introduced a greedy heuristic that achieves strong empirical results.
@@ 1.2 Our Results
  • For the Chromatic Correlation Clustering problem, we enhance the LP-based method through a new analysis that yields a $2.15$-approximation, improving upon the previous best bound of $2.5$ by Xiu et al. XiuHTCH22 . We also establish a lower bound of $2.11$ within the same analytical framework, underscoring the near-optimality of our approach.
@@ 2 Preliminaries
The correlation clustering (CC) problem takes as input a signed undirected graph $G=(V,E=E^{+}\uplus E^{-})$, where each edge $e=uv\in E$ is assigned a sign ‘++’ or ‘-’, described by $e\in E^{+}$ or $e\in E^{-}$. The objective is to find a partition of the nodes such that the number of disagreements—i.e., negative edges within the same cluster and positive edges between different clusters—is minimized. In other words, the cost of the clustering $\mathcal{C}$ is as follows:
@@ 2.1 Pseudometric-weighted Correlation Clustering
The weighted Correlation Clustering problem is a generalization of the classical CC problem in which each edge is associated with a nonnegative violation cost. Specifically, for each edge $uv$, a weight $w_{uv}\geq 0$ is provided, and violating the edge’s label (either ‘++’ or ‘-’) incurs a penalty of $w_{uv}$. This differs from the standard setting, where all violations incur a uniform cost of $1$. The weighted variant allows us to encode edge-wise reliability: when $w_{uv}$ is large, it is more reasonable to follow the given label between $u$ and $v$.
However, assuming the Unique Games Conjecture, no $O(1)$-approximation algorithm exists for the general weighted case DEMAINE2006172 . An exception occurs when the weight function satisfies the triangle inequality, i.e., the weights form a pseudometric. In this pseudometric-weighted setting, a constant-factor approximation is known DBLP:conf/soda/CharikarG24 . Following the analysis of Charikar and Gao with $L=2$ yields an approximation factor of $\overline{B}_{\text{HR}}+\frac{1}{\frac{1}{3}}\leq\frac{4}{\frac{1}{3}}+2(L-1)+\frac{1}{\frac{1}{3}}=17,$ since the second type of charge occurs at most L1=1L-1=1 time in the charging scheme. The following is a natural LP relaxation of the weighted CC problem, extending (CC-LP):
@@ 2.2 Chromatic Correlation Clustering Problem
The Chromatic Correlation Clustering problem is a variant of the classical CC problem in which each cluster is additionally assigned a color bonchi2012chromatic . The input includes a set of $L$ possible colors, as well as a special color $\gamma$ that denotes that two vertices should not be placed in the same cluster—analogous to a negative (‘-’) edge in the classical CC setting. When $L=1$ (i.e., a single cluster color), CCC reduces to the standard CC problem.
@@
Since CCC generalizes the standard CC problem, the integrality gap of (CCC-LP) is at least as large as that of ($C$-LP), which is 2.
@@ 3 Approximation Algorithm
Input: Graph G=(V,E)G=(V,\,E), color function $\phi:E\rightarrow L\cup\{\gamma\}$, LP solution $\{x_{u}^{c}\}_{u\in V,\,c\in L}$ and $\{x_{uv}^{c}\}_{uv\in E,\,c\in L}$.
@@
  if $\exists c\in L$ s.t. $x_{u}^{c}<\frac{1}{2}$, then
@@
[…]
@@ 4 Rounding Functions
The effectiveness of the LP-Pivot and LP-CCC algorithms critically depends on the choice of rounding functions used in the clustering process. Rounding functions $f^{+},\,f^{-},\,f^{\circ}:[0,1]\rightarrow[0,1]$ convert the LP value $x_{uv}$ to the non-selection probability $p_{uv}$ chawla2015near . The sign of the edge uvuv—either ‘++’, ‘-’, or $\circ$’—determines which rounding function is applied. The sign $\circ$’ indicates that the edge does not belong to $E$. The following natural conditions are imposed on any rounding function $f$:
@@ 4.3 Comparison with Prior Work
Pseudometric-weighted CC: The LP-UMVD-Pivot algorithm, recently proposed by Charikar and Gao DBLP:conf/soda/CharikarG24 for the Ultrametric Violation Distance (UMVD) problem, follows a pivoting-based rounding strategy applied to an LP relaxation. When the number of distinct pairwise distances between elements, denoted by $L$, is equal to 2, their algorithm can be viewed as a special case of our LP-Pivot framework. In this setting, the two distances $d_{1}$ and $d_{2}$ correspond to the ‘-’ and ‘++’ labels, respectively, used in our rounding procedure. The rounding functions used are:
$f^{+}(x)=f^{-}(x)=\begin{cases}0,&x<\alpha;\\ \frac{\max\{x-\alpha\beta,\,0\}}{1-\alpha\beta},&\alpha\leq x\leq 1-\alpha;\\ 1,&x>1-\alpha,\end{cases}\qquad f^{\circ}(x)=\begin{cases}0,&x<\alpha\beta;\\ x,&\alpha\beta\leq x\leq 1-\alpha\beta;\\ 1,&x>1-\alpha\beta.\end{cases}$
@@
With triple-based analysis, this choice yields an approximation factor of $6$, as shown in the subsection 4.4.
@@
4.4 Proof for $6$-approximation on Pseudometric-weighted CC with LP-UMVD-Pivot
Since the formulas for $e.cost$ and $e.lp$ in the pseudometric-weighted CC are compatible with those in the classical CC, and weights $w$ affect $\mathcal{C}$ linearly, the following lemma presented by Chawla et al. chawla2015near remains applicable.
Lemma 2 (Lemma 5 of chawla2015near ).
  1. 1.
    the triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$; or
  2. 2.
We introduce another useful lemma:
Lemma 3.
Proof.
$\alpha\cdot e.lp_{w}-e.cost_{w}$ $\geq\alpha(1-p_{vw}p_{wu})\min\{x,1-x\}$
$\quad-\max\{p_{vw}(1-p_{wu})+(1-p_{vw})p_{wu},(1-p_{vw})(1-p_{wu})\}$
$\geq\alpha(1-p_{vw}p_{wu})\cdot\frac{1}{\alpha}-(1-p_{vw}p_{wu})=0.$
4.4.1 Triangle inequality is tight.
Figure 2: Configuration of regions induced from (14).
Table 1: Formula for $e.cost$ and $e.lp$ for each configuration.
$x_{uv}=x$ $x_{vw}=y$ $x_{wu}=z=x+y$
Region Sign $e.cost_{w}$ $e.lp_{w}$ $e.cost_{u}$ $e.lp_{u}$ $e.cost_{v}$ $e.lp_{v}$
I ++ 0 $x$ 0 $y$ 0 $z$
- 11 1x1-x 11 1y1-y 11 1z1-z
II ++ $z$ $x$ $z$ $y$ 0 $z$
- 1z1-z 1x1-x 1z1-z 1y1-y 11 1z1-z
III ++ y+z2yzy+z-2yz (1yz)x(1-yz)x $z$ $y$ $y$ $z$
- (1y)(1z)(1-y)(1-z) (1yz)(1x)(1-yz)(1-x) 1z1-z 1y1-y 1y1-y 1z1-z
IV ++ 1y1-y (1y)x(1-y)x 11 $y$ $y$ $z$
- 0 (1y)(1x)(1-y)(1-x) 0 1y1-y 1y1-y 1z1-z
V ++ 0 0 11 $y$ 11 $z$
- 0 0 0 1y1-y 0 1z1-z
VI ++ 1y1-y (1y)x(1-y)x 1x1-x (1x)y(1-x)y x+y2xyx+y-2xy (1xy)z(1-xy)z
- 0 (1y)(1x)(1-y)(1-x) 0 (1x)(1y)(1-x)(1-y) (1x)(1y)(1-x)(1-y) (1xy)(1z)(1-xy)(1-z)
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Region I, II, V
I $(0,0)$ ++ 0 0 0
- 55 55 55
$(0,\frac{1}{3})$ ++ 0 22 22
- 55 33 33
$(\frac{1}{6},\frac{1}{6})$ ++ 11 11 22
- 44 44 33
II $(0,\frac{1}{3})$ ++ 1/3-1/3 $5/3$ 22
- $16/3$ $10/3$ 33
$(\frac{1}{6},\frac{1}{6})$ ++ $2/3$ $2/3$ 22
- $13/3$ $1/3$ 33
$(\frac{1}{3},\frac{1}{3})$ ++ $4/3$ $4/3$ 44
- $11/3$ $11/3$ 11
V $(0,\frac{2}{3})$ ++ 0 33 33
- 0 22 22
$(0,1)$ ++ 0 55 55
- 0 0 0
$(\frac{1}{3},\frac{2}{3})$ ++ 0 33 55
- 0 22 0
Region III
$y+z-2yz-(1-y)(1-z)=\frac{1-(2-3y)(2-3z)}{3}\geq 0$
since $y,z\in[1/3,2/3]$ within the region. Therefore, $6\cdot e.lp_{w}-e.cost_{w}$ is smaller with the sign ‘++’.
Now, consider the minimum value possible for the formula above within the region. Replacing x=zyx=z-y,
$6\cdot e.lp_{w}-e.cost_{w}=-6yz^{2}+6y^{2}z+2yz-7y+5z.$
Fixing $y$, the formula is concave of $z$; therefore, it is minimized in $z=y$ or $z=2/3$, resulting in $2y^{2}-2y$ and $4y^{2}-\frac{25}{3}y+\frac{10}{3}$ each. Within the range $y\in[1/3,2/3]$, the minimum value of each is 1/2-1/2 in $y=1/2$ and 4/9-4/9 in $y=2/3$. Thus, $6\cdot e.lp_{w}-e.cost_{w}\geq-1/2$.
Finally, consider the minimum value possible for $6\cdot e.lp_{u}-e.cost_{u}$ and $6\cdot e.lp_{v}-e.cost_{v}$. Again, these are affine; hence checking on extremal points is sufficient. It turns out that $6\cdot e.lp_{u}-e.cost_{u}\geq 4/3$, minimized at $(y,z)=(1/3,2/3)$ with ‘++’ sign; $6\cdot e.lp_{v}-e.cost_{v}\geq 4/3$, minimized at $(y,z)=(1/3,2/3)$ with ‘-’ sign.
Region IV
We can utilize the optimization method explained in 4.4 again: Since $(1-y)x\leq(1-y)(1-x)$ and $1-y\geq 0$, the formula with the sign ‘++’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is (6x1)(1y)(6x-1)(1-y). This has a minimum value of 1/3-1/3 in $(x,y)=(0,2/3)$.
Also, since $6z-y=6x+5y\geq 10/3>5/2$ in this region, the formula with the sign ‘-’ is smaller for $6\cdot e.lp_{v}-e.cost_{v}$, which is 56x5y5-6x-5y. This has a minimum value of 1/3-1/3 in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is 6y16y-1 or 6(1y)6(1-y). Each of them has a minimum value $1$ in $(x,y)=(1/3,1/3)$, 22 in $(x,y)=(0,2/3)$, respectively. Thus, $(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{u}-e.cost_{u})$ and $(6\cdot e.lp_{u}-e.cost_{u})+(6\cdot e.lp_{v}-e.cost_{v})$ are at least $1/3+1=2/3\geq 0$.
Finally,
$=-6xy-4y+4\geq-6y+4\geq 0,$
Region VI
As in IV, the formula with sign ‘++’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is (6x1)(1y)(6x-1)(1-y). This has a minimum value of $1/3$ in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is (1x)(6y1)(1-x)(6y-1) or 6(1x)(1y)6(1-x)(1-y). Each of them has a minimum value $2/3$ in $(x,y)=(1/3,1/3)$, $4/3$ in $(x,y)=(1/3,2/3)$, respectively.
For $6\cdot e.lp_{v}-e.cost_{v}$, consider the difference between those of ‘++’ sign and ‘-’ sign.
$=xy-2x-2y+3$
$=(2-x)(2-y)-1\geq\frac{11}{9}>0,$
since $z\geq 2/3$. Therefore, the formula with sign ‘-’ is smaller, which is 6(1xy)(1z)(1x)(1y)6(1-xy)(1-z)-(1-x)(1-y).
Define $p=xy$, then
6(1xy)(1z)(1x)(1y)=6pz5z7p+5,6(1-xy)(1-z)-(1-x)(1-y)=6pz-5z-7p+5,
with region $z\in[2/3,1],\,p\in\left[\frac{z}{3}-\frac{1}{9},\left(\frac{z}{2}\right)^{2}\right]$. Fixing the value of $z$, it is a decreasing function with respect to $p$; fixing the value of $p$, it is a decreasing function with respect to $z$. Therefore, the function is minimized at $z=1,\,p=\frac{1}{4}$ with the value $\frac{1}{4}$.
Overall, $6\cdot LP-ALG\geq 0$ when the triangle inequality is tight.
4.4.2 All coordinates belong to endpoint of segments.
There are total $4$ possible LP values: 0, 1/3, 2/3, 10,\,1/3,\,2/3,\,1. Among all the possible $(x,y,z)$ tuples, there are only $7$ tuples whose triangle inequality is not tight. Moreover, $3$ tuples $(x,y,z)=(\frac{1}{3},\frac{1}{3},\frac{1}{3}),\,(\frac{1}{3},\frac{2}{3},\frac{2}{3}),\,(\frac{2}{3},\frac{2}{3},\frac{2}{3})$ among them only consist of value $1/3$ or $2/3$, thus $3\cdot LP-ALG\geq 0$ by Lemma 3. Therefore, we only need to verify the remaining $3$ tuples.
Since left-side and right-side limits of the value of rounding function at such point might differ, we have to consider all possible limits of $6\cdot LP-ALG$ up to $2^{3}=8$ directions. Table 3 shows remaining $3$ tuples and the directional limit of the value of $6\cdot e.cost-e.lp$ for each configuration. Note that since $x\leq y\leq z$, only up to $4$ out of $8$ directions are required to verification.
$(\frac{2}{3}-\delta,\frac{2}{3}-\delta,1)$ ++ 11 11 $26/9$
- $2/3$ $2/3$ 1/9-1/9
$(\frac{2}{3}-\delta,\frac{2}{3}+\delta,1)$ ++ 0 11 $5/3$
- 0 $2/3$ 0
$(\frac{2}{3}+\delta,\frac{2}{3}+\delta,1)$ ++ 0 0 0
- 0 0 0
$(\frac{1}{3}-\delta,1,1)$ ++ 0 55 55
- 0 0 0
$(\frac{1}{3}+\delta,1,1)$ ++ 0 $10/3$ $10/3$
- 0 0 0
$(\frac{2}{3}-\delta,1,1)$ ++ 0 $5/3$ $5/3$
- 0 0 0
$(\frac{2}{3}+\delta,1,1)$ ++ 0 0 0
- 0 0 0
$(1,1,1)$ ++ 0 0 0
- 0 0 0
This completes our case analysis.